# Preconditioning without a preconditioner using randomized block Krylov subspace methods

### v2

@@ 2.2.2 Nyström preconditioning
Techniques from randomized numerical linear algebra allow near-optimal low-rank approximations to $\mathbf{A}$ to be computed very efficiently [34, 61]. Intuitively, we might hope that the corresponding preconditioner works nearly as well as spectral deflation.
@@
where $\mathbf{K}_{s}:=[\mathbf{\Omega}~\mathbf{A}\mathbf{\Omega}~\cdots~\mathbf{A}^{s-1}\mathbf{\Omega}]\in\mathbb{R}^{d\times(s\ell)}$ and $\mathbf{\Omega}\in\mathbb{R}^{d\times\ell}$ is a matrix of independent standard normal random variables. [^3] In this paper, we only consider Nyström preconditioning where $\mathbf{\Omega}$ is a Gaussian matrix. Variants of Nyström approximation based on subsampling rows/columns of $\mathbf{A}$ are effective in some settings (e.g. Kernel Ridge Regression); see [23, §2.2.3] for a discussion. This variant of the Nyström approximation is among the most powerful randomized low-rank approximation algorithms, and can be implemented using $s$ matrix-loads. Note, however, that the algorithm should not be implemented as written in Eq. 12. In particular, one should avoid using a monomial basis $\mathbf{K}_{s}$ and carefully structure the interactions with $\mathbf{A}$ to avoid unnecessary costs; see e.g. [61] for pseudocode.
Nyström-based preconditioning is effective in theory and practice, and has been an active area of research in recent years [40, 23, 5, 20, 65, 36, 15]. Most related to the content of this paper is the theoretical analysis of [23], which proves that if $s=1$, $\theta=\lambda_{\ell}(\mathbf{A}\langle\mathbf{\Omega}\rangle)$, and the sketching dimension $\ell$ is on the order of the effective dimension $d_{\textup{eff}}(\mu):=\operatorname{tr}\bigl{(}\mathbf{A}\mathbf{A}_{\mu}^{-1}\bigr{)}=\sum_{i=1}^{d}{\lambda_{i}}/{(\lambda_{i}+\mu)}$, then Nyström PCG converges in at most $t=O(\log(1/\varepsilon))$ iterations; i.e. independent of any spectral properties of $\mathbf{A}$. Our analysis makes use of the same general techniques as [23], but is applicable when $s>1$ as well as if $\mu=0$. We discuss the theoretical bounds of [23] in Section B.2.
@@ 3 Our approach: augmented block-CG
In this paper, we advocate for the direct application of block KSMs to Eq. 1. The concept of block Krylov subspaces Eq. 6 naturally gives rise to the block-CG algorithm [48].
@@
$\mathsf{bcg}_{t}^{(i)}(\mu):=\operatornamewithlimits{argmin}_{\mathbf{x}\in\mathcal{K}_{t}(\mathbf{A},\mathbf{B})}\|\mathbf{A}_{\mu}^{-1}\mathbf{b}^{(i)}-\mathbf{x}\|_{\mathbf{A}_{\mu}}.$
The block-CG iterates $\mathsf{bcg}_{t}^{(1)}(\mu),\ldots,\mathsf{bcg}_{t}^{(m)}(\mu)$ can be simultaneously computed using t1t-1 block matrix-vector products with $\mathbf{A}$. We discuss implementation and costs further in Section 3.2.
@@ 3.1 Implicit preconditioning
Our main conceptual result is that by augmenting $\mathbf{b}$ with $\mathbf{\Omega}$, block-CG implicitly enjoys the benefits of certain classes of preconditioners built using $\mathbf{\Omega}$. We begin with a key observation about the relation between the block Krylov subspace Eq. 6 and the preconditioned Krylov subspace Eq. 8 corresponding to a certain class of preconditioners.
@@
(13) $(\mathbf{P}_{\mu}^{-1}\mathbf{A})^{k}\mathbf{P}_{\mu}^{-1}\mathbf{b}=((\mathbf{I}+\mathbf{X})\mathbf{A})^{k}(\mathbf{I}+\mathbf{X})\mathbf{b},\quad k=0,1,\ldots,t-1,$
and each $((\mathbf{I}+\mathbf{X})\mathbf{A})^{k}(\mathbf{I}+\mathbf{X})\mathbf{b}$ can be expressed as linear combination of vectors which live in the specified space.
@@
Since (block) Krylov subspaces are shift invariant, without loss of generality it suffices to consider the case $\mu=0$. For notational simplicity we will denote $\mathbf{P}_{\mu}$ by $\mathbf{P}$.
@@
(14) $\mathcal{K}_{1}(\mathbf{P}^{-1}\mathbf{A},\mathbf{P}^{-1}\mathbf{b})=\operatorname{span}(\mathbf{P}^{-1}\mathbf{b})=\operatorname{span}((\mathbf{I}+\mathbf{X})\mathbf{b})=\operatorname{span}(\mathbf{b}+\mathbf{X}\mathbf{b}).$
Clearly $\mathbf{b}\in\mathcal{K}_{1}(\mathbf{A},\mathbf{b})$ and $\mathbf{X}\mathbf{b}\in\operatorname{range}(\mathbf{X})\subseteq\mathcal{K}_{s+1}(\mathbf{A},\mathbf{\Omega})$, so $\mathbf{\mathbf{b}+\mathbf{X}\mathbf{b}}\subseteq\mathcal{K}_{1}(\mathbf{A},\mathbf{b})+\mathcal{K}_{s+t}(\mathbf{A},\mathbf{\Omega})$ as desired.
@@
(15) $\mathcal{K}_{t-1}(\mathbf{P}^{-1}\mathbf{A},\mathbf{P}^{-1}\mathbf{b})\subseteq\mathcal{K}_{t-1}(\mathbf{A},\mathbf{b})+\mathcal{K}_{(t-1)+s}(\mathbf{A},\mathbf{\Omega}).$
@@
(16) $\hskip 130.0002pt(\mathbf{P}^{-1}\mathbf{A})^{t-2}\mathbf{P}^{-1}\mathbf{b},(\mathbf{P}^{-1}\mathbf{A})^{t-1}\mathbf{P}^{-1}\mathbf{b}\}.$
@@
(17) $(\mathbf{P}^{-1}\mathbf{A})^{j}\mathbf{P}^{-1}\mathbf{b}$ $\in\mathcal{K}_{t-1}(\mathbf{P}^{-1}\mathbf{A},\mathbf{P}^{-1}\mathbf{b})$
(18) $\subseteq\mathcal{K}_{t-1}(\mathbf{A},\mathbf{b})+\mathcal{K}_{t+s-1}(\mathbf{A},\mathbf{\Omega})$
(19) $\subseteq\mathcal{K}_{t}(\mathbf{A},\mathbf{b})+\mathcal{K}_{t+s}(\mathbf{A},\mathbf{\Omega}),$
@@
(20) $(\mathbf{P}^{-1}\mathbf{A})^{t-1}\mathbf{P}^{-1}\mathbf{b}=(\mathbf{P}^{-1}\mathbf{A})\mathbf{v}=(\mathbf{I}+\mathbf{X})\mathbf{A}\mathbf{v}=\mathbf{A}\mathbf{v}+\mathbf{X}\mathbf{A}\mathbf{v}.$
@@
(21) $\mathbf{A}\mathbf{v}\in\mathbf{A}\mathcal{K}_{t-1}(\mathbf{A},\mathbf{b})+\mathbf{A}\mathcal{K}_{t+s-1}(\mathbf{A},\mathbf{\Omega})\subseteq\mathcal{K}_{t}(\mathbf{A},\mathbf{b})+\mathcal{K}_{t+s}(\mathbf{A},\mathbf{\Omega}).$
@@
Theorem 3.2 allows the performance of augmented block-CG to be related to PCG with a broad choice of preconditioner.
@@
Fix any matrix $\mathbf{\Omega}\in\mathbb{R}^{d\times m}$ and let $\mathbf{P}=(\mathbf{I}+\mathbf{X})^{-1}$ be any preconditioner where $\operatorname{range}(\mathbf{X})\subseteq\mathcal{K}_{s+1}(\mathbf{A},\mathbf{\Omega})$. Define the augmented starting block $\mathbf{B}=[\mathbf{b}~\mathbf{\Omega}]$. Then, for any $t\geq s$, the $t$-th block-CG iterate is related to the (ts)(t-s)-th preconditioned-CG iterate corresponding to the preconditioner $\mathbf{P}_{\mu}$ in that
@@
Clearly $\mathcal{K}_{t}(\mathbf{A},\mathbf{b})+\mathcal{K}_{t+s}(\mathbf{A},\mathbf{\Omega})\subseteq\mathcal{K}_{t+s}(\mathbf{A},[\mathbf{b}~\mathbf{\Omega}])$, so the result follows immediately from Theorem 3.2, the optimality of block-CG, and the definition of preconditioned-CG.
@@
Theorem 3.4 asserts that augmented block-CG automatically performs no worse than Nyström PCG (with the best choice of $s$ and $\theta$) after the same number of matrix-loads. [^4] Recall, $s$ matrix-loads are used to compute the Nyström approximation $\mathbf{A}\langle\mathbf{K}_{s}\rangle$, whose range lives in $\mathcal{K}_{s+1}(\mathbf{A},\mathbf{\Omega})$. In particular, when $\mathbf{U}\mathbf{D}\mathbf{U}^{\mathsf{T}}=\mathbf{A}\langle\mathbf{K}_{s}\rangle$, then the deflation preconditioner $\mathbf{P}_{\mu}$ defined in Eq. 9 has the form
(22) $\mathbf{P}^{-1}=\mathbf{I}+\mathbf{X},\quad\text{where $\operatorname{range}(\mathbf{X})\subseteq\mathcal{K}_{s+1}(\mathbf{A},\mathbf{\Omega})$}.$
This explains the performance observed in Fig. 1.
In Section 4, we use use Theorem 3.4, and a new bound for Nyström preconditioning (Theorem 4.7), to prove a bound for block-CG reminiscent of Corollary 2.4 for the spectral deflation preconditioner.
@@ 3.2 Computational costs
(23) $\operatorname{range}(\mathbf{Q}_{t})=\mathcal{K}_{t}(\mathbf{A},\mathbf{B}),\quad\mathbf{T}_{t}=\mathbf{Q}_{t}^{\mathsf{T}}\mathbf{A}\mathbf{Q}_{t}.$
@@
where $\mathbf{e}_{1}=[1,0,\ldots,0]^{\mathsf{T}}$. In particular, since $\mathcal{K}_{t}(\mathbf{A}_{\mu},\mathbf{B})=\mathcal{K}_{t}(\mathbf{A},\mathbf{B})$ for any scalar $\mu$, Eq. 24 can be used to compute the block-CG iterate for multiple values of $\mu$ using the same $\mathbf{Q}_{t}$ and $\mathbf{T}_{t}$. For more details on block-Lanczos and the connection to block-CG, see [48, 55, 7, 60, etc.].
There are many variants of block-Lanczos, since there are many possible orthogonlization schemes that can be use within block-Lanczos type algorithms [6, 47, 1]. The following provides a high-level overview of the costs associated with block-Lanczos with full-reorthogonalization.
@@
  • $O(dm^{2}t^{2})$ floating point operations, and
  • $O(dmt)$ storage.
@@
Proof 3.8 (Proof of Theorem 3.7).
At each iteration the block-Lanczos algorithm performs one matrix-load ($m$ parallel matrix products) and lower-order arithmetic. The dominant storage cost is storing $\mathbf{Q}_{t}$, which has $mt$ columns. The arithmetic cost is dominated by the $O(dm^{2}t^{2})$ operations required to obtain $\mathbf{Q}_{t}$ using full reorthogonalization. Since $\mathbf{T}_{t}$ (and hence $\mathbf{T}_{t}+\mu\mathbf{I}$) are $(mt)\times(mt)$ matrices of bandwidth $O(m)$, the linear system $(\mathbf{T}_{t}+\mu\mathbf{I})\mathbf{z}=\mathbf{e}_{1}$ can be solved in $O((mt)m^{2})=O(m^{3}t)=O(dm^{2})$ time. Subsequently, since $\mathbf{Q}_{t}$ is a $d\times(mt)$ matrix, $\mathbf{Q}_{t}\mathbf{z}$ can be computed in $O(dmt)$ time.
We emphasize that after the block Lanczos algorithm has been run, the cost to compute the block-CG iterates for multiple values of $\mu$ is relatively small. This is in contrast to Nyström PCG, which is is not particularly well suited for solving Eq. 1 for multiple values of $\mu$; while the Nyström approximation can be reused, PCG must be re-run and new products with $\mathbf{A}$ computed.
Notice that the costs (besides matrix-loads) scale with the block-size $m$. We again remind the reader that, while Theorem 3.4 guarantees augmented block-CG outperforms Nyström preconditioning in terms of matrix-loads, but our theory provides no guarantees for other cost matrices like matrix-products or floating point operations. Users should carefully consider the costs and benefits associated with the various algorithms in the context of their own computing environments.
@@
By avoiding orthogonalizing and storing the whole Krylov basis, it is in fact possible to reduce the floating point cost to $O(dmtL+m^{3}tL)$ and the storage to $O(dm+dL)$, where $L$ is the number of values of $\mu$ at which one wishes to evaluate $\mathsf{bcg}_{t}^{(1)}(\mu)$ [48]. However, there are several caveats to such an implementation. First, such methods require knowing the values of $\mu$ at which one wishes to evaluate $\mathsf{bcg}_{t}^{(1)}(\mu)$ ahead of time. Second, the finite-precision behavior of such methods can differ greatly from the exact arithmetic behavior. We explore the impacts of finite precision arithmetic Appendix A, where we perform some numerical experiments that indicate that full reorthogonalization may be unnecessary in some situations.
For methods based on the standard Lanczos algorithm, there is quite a bit of theory which guarantees that such methods can still work, even without orthogonalization [49, 50, 19, 28, 46, 7]. Unfortunately, much less theory is known about the block-Lanczos algorithm in finite precision arithmetic [67]. Without reorthogonalization, we observe that in many cases the block-CG method fails to converge at all, while in other cases it does converge for some time; see Appendix A. Understanding this behavior is well beyond the scope of the current work.

### v3

@@ 2.2.2 Nyström preconditioning
Techniques from randomized numerical linear algebra allow near-optimal low-rank approximations to $\mathbf{A}$ to be computed very efficiently [38, 68], so we might hope that the corresponding preconditioner works nearly as well as spectral deflation, while being much cheaper to compute.
@@
where $\mathbf{K}_{s}:=[\mathbf{\Omega}~\mathbf{A}\mathbf{\Omega}~\cdots~\mathbf{A}^{s-1}\mathbf{\Omega}]\in\mathbb{R}^{d\times(s\ell)}$ and $\mathbf{\Omega}\in\mathbb{R}^{d\times\ell}$ is a matrix of independent standard normal random variables.§§§In this paper, we only consider Nyström preconditioning where $\mathbf{\Omega}$ is a Gaussian matrix. Variants of Nyström approximation based on subsampling rows/columns of $\mathbf{A}$ are effective in some settings (e.g. Kernel Ridge Regression); see [25, §2.2.3] for a discussion. This variant of the Nyström approximation is among the most powerful randomized low-rank approximation algorithms, and can be implemented using $s$ matrix-loads. Note, however, that the algorithm should not be implemented as written in Equation˜9. In particular, one should avoid using a monomial basis $\mathbf{K}_{s}$ and carefully structure the interactions with $\mathbf{A}$ to avoid unnecessary costs; see e.g. [68] for pseudocode.
Nyström-based preconditioning is effective in theory and practice, and has been an active area of research in recent years [46, 25, 6, 21, 73, 41, 16]. Most related to the content of this paper is the theoretical analysis of [25], which proves that if $s=1$, $\theta=\lambda_{\ell}(\mathbf{A}\langle\mathbf{\Omega}\rangle)$, and the sketching dimension $\ell$ is on the order of the effective dimension $d_{\textup{eff}}(\mu):=\operatorname{tr}\bigl(\mathbf{A}\mathbf{A}_{\mu}^{-1}\bigr)=\sum_{i=1}^{d}{\lambda_{i}}/{(\lambda_{i}+\mu)}$, then Nyström PCG converges in at most $t=O(\log(1/\varepsilon))$ iterations; i.e. independent of any spectral properties of $\mathbf{A}$. Our analysis makes use of the same general techniques as [25], but is applicable when $s>1$ as well as if $\mu=0$. We discuss the theoretical bounds of [25] in Subsection˜B.2.
@@ 3 Our approach: augmented block-CG
In this paper, we advocate for the direct application of block KSMs to Equation˜1. The concept of block Krylov subspaces Equation˜6 naturally gives rise to the block-CG algorithm [54].
@@
$\mathsf{bcg}_{t}^{(i)}(\mu):=\operatornamewithlimits{argmin}_{\mathbf{x}\in\mathcal{K}_{t}(\mathbf{A}_{\mu},\mathbf{B})}\|\mathbf{A}_{\mu}^{-1}\mathbf{b}^{(i)}-\mathbf{x}\|_{\mathbf{A}_{\mu}}.$
The block-CG iterates $\mathsf{bcg}_{t}^{(1)}(\mu),\ldots,\mathsf{bcg}_{t}^{(m)}(\mu)$ can be simultaneously computed using t1t-1 block matrix-vector products with $\mathbf{A}$. We discuss implementation and costs further in Subsection˜3.2.
@@ 3.1 Implicit preconditioning
Our main conceptual result is that by augmenting $\mathbf{b}$ with $\mathbf{\Omega}$, block-CG implicitly enjoys the benefits of certain classes of preconditioners built using $\mathbf{\Omega}$. We begin with a key observation about the relation between the block Krylov subspace Equation˜6 and the preconditioned Krylov subspace Equation˜7 corresponding to a certain class of preconditioners.
@@
$(\mathbf{P}_{\mu}^{-1}\mathbf{A}_{\mu})^{k}\mathbf{P}_{\mu}^{-1}\mathbf{b}=((\mathbf{I}+\mathbf{X})\mathbf{A}_{\mu})^{k}(\mathbf{I}+\mathbf{X})\mathbf{b},\quad k=0,1,\ldots,t-1,$
and each $((\mathbf{I}+\mathbf{X})\mathbf{A}_{\mu})^{k}(\mathbf{I}+\mathbf{X})\mathbf{b}$ can be expressed as linear combination of vectors which live in the specified space.
@@
By the shift invariance of (block) Krylov subspaces, $\mathcal{K}_{t}(\mathbf{A},\mathbf{b})+\mathcal{K}_{t+s}(\mathbf{A},{\Omega})=\mathcal{K}_{t}(\mathbf{A}_{\mu},\mathbf{b})+\mathcal{K}_{t+s}(\mathbf{A}_{\mu},\mathbf{\Omega})$. Hence, without loss of generality it suffices to consider the case $\mu=0$. For notational simplicity we will denote $\mathbf{P}_{\mu}$ by $\mathbf{P}$.
@@
$\mathcal{K}_{1}(\mathbf{P}^{-1}\mathbf{A},\mathbf{P}^{-1}\mathbf{b})=\operatorname{span}(\mathbf{P}^{-1}\mathbf{b})=\operatorname{span}((\mathbf{I}+\mathbf{X})\mathbf{b})=\operatorname{span}(\mathbf{b}+\mathbf{X}\mathbf{b}).$
Clearly $\mathbf{b}\in\mathcal{K}_{1}(\mathbf{A},\mathbf{b})$ and $\mathbf{X}\mathbf{b}\in\operatorname{range}(\mathbf{X})\subseteq\mathcal{K}_{s+1}(\mathbf{A},\mathbf{\Omega})$, so $\mathbf{\mathbf{b}+\mathbf{X}\mathbf{b}}\subseteq\mathcal{K}_{t}(\mathbf{A},\mathbf{b})+\mathcal{K}_{t+s}(\mathbf{A},\mathbf{\Omega})$ as desired.
@@
$\mathcal{K}_{t-1}(\mathbf{P}^{-1}\mathbf{A},\mathbf{P}^{-1}\mathbf{b})\subseteq\mathcal{K}_{t-1}(\mathbf{A},\mathbf{b})+\mathcal{K}_{(t-1)+s}(\mathbf{A},\mathbf{\Omega}).$
@@
$\hskip 130.0002pt(\mathbf{P}^{-1}\mathbf{A})^{t-2}\mathbf{P}^{-1}\mathbf{b},(\mathbf{P}^{-1}\mathbf{A})^{t-1}\mathbf{P}^{-1}\mathbf{b}\}.$
@@
$(\mathbf{P}^{-1}\mathbf{A})^{j}\mathbf{P}^{-1}\mathbf{b}$ $\in\mathcal{K}_{t-1}(\mathbf{P}^{-1}\mathbf{A},\mathbf{P}^{-1}\mathbf{b})$
$\subseteq\mathcal{K}_{t-1}(\mathbf{A},\mathbf{b})+\mathcal{K}_{t+s-1}(\mathbf{A},\mathbf{\Omega})$
$\subseteq\mathcal{K}_{t}(\mathbf{A},\mathbf{b})+\mathcal{K}_{t+s}(\mathbf{A},\mathbf{\Omega}),$
@@
$(\mathbf{P}^{-1}\mathbf{A})^{t-1}\mathbf{P}^{-1}\mathbf{b}=(\mathbf{P}^{-1}\mathbf{A})\mathbf{v}=(\mathbf{I}+\mathbf{X})\mathbf{A}\mathbf{v}=\mathbf{A}\mathbf{v}+\mathbf{X}\mathbf{A}\mathbf{v}.$
@@
$\mathbf{A}\mathbf{v}\in\mathbf{A}\mathcal{K}_{t-1}(\mathbf{A},\mathbf{b})+\mathbf{A}\mathcal{K}_{t+s-1}(\mathbf{A},\mathbf{\Omega})\subseteq\mathcal{K}_{t}(\mathbf{A},\mathbf{b})+\mathcal{K}_{t+s}(\mathbf{A},\mathbf{\Omega}).$
@@
Theorem˜3.2 allows the performance of augmented block-CG to be related to PCG with a broad choice of preconditioner.
@@
Fix any matrix $\mathbf{\Omega}\in\mathbb{R}^{d\times m}$ and let $\mathbf{P}_{\mu}=(\mathbf{I}+\mathbf{X})^{-1}$ be any preconditioner where $\operatorname{range}(\mathbf{X})\subseteq\mathcal{K}_{s+1}(\mathbf{A},\mathbf{\Omega})$. Define the augmented starting block $\mathbf{B}=[\mathbf{b}~\mathbf{\Omega}]$. Then, for any $t\geq s$, the $t$-th block-CG iterate is related to the (ts)(t-s)-th preconditioned-CG iterate corresponding to the preconditioner $\mathbf{P}_{\mu}$ in that
@@
Clearly $\mathcal{K}_{t}(\mathbf{A},\mathbf{b})+\mathcal{K}_{t+s}(\mathbf{A},\mathbf{\Omega})\subseteq\mathcal{K}_{t+s}(\mathbf{A},[\mathbf{b}~\mathbf{\Omega}])$, so the result follows immediately from Theorem˜3.2, the optimality of block-CG, and the definition of preconditioned-CG.
@@
Theorem˜3.4 asserts that augmented block-CG automatically performs no worse than Nyström PCG (with the best choice of $s$ and $\theta$) after the same number of matrix-loads.Recall, $s$ matrix-loads are used to compute the Nyström approximation $\mathbf{A}\langle\mathbf{K}_{s}\rangle$, whose range lives in $\mathcal{K}_{s+1}(\mathbf{A},\mathbf{\Omega})$. In particular, when $\mathbf{U}\mathbf{D}\mathbf{U}^{\mathsf{T}}=\mathbf{A}\langle\mathbf{K}_{s}\rangle$,where $\langle\mathbf{K}_{s}\rangle$ is the Nyström approximation Equation˜9, then the deflation preconditioner $\mathbf{P}_{\mu}$ defined in Equation˜8 has the form
$\mathbf{P}^{-1}=\mathbf{I}+\mathbf{X},\quad\text{where $\operatorname{range}(\mathbf{X})\subseteq\mathcal{K}_{s+1}(\mathbf{A},\mathbf{\Omega})$}.$
This is a deterministic statement that explains the relative performance of the algorithms in Figure˜1.
In Section˜4, we use use Theorem˜3.4, and a new bound for Nyström preconditioning (Theorem˜4.7) when $\mathbf{\Omega}$ is a Gaussian random matrix to prove probabilistic bounds for block-CG reminiscent of Corollary˜2.4 for the spectral deflation preconditioner.
@@ 3.2 Computational costs
$\operatorname{range}(\mathbf{Q}_{t})=\mathcal{K}_{t}(\mathbf{A},\mathbf{B}),\quad\mathbf{T}_{t}=\mathbf{Q}_{t}^{\mathsf{T}}\mathbf{A}\mathbf{Q}_{t}.$
@@
where $\mathbf{e}_{1}=[1,0,\ldots,0]^{\mathsf{T}}$. In particular, since $\mathcal{K}_{t}(\mathbf{A}_{\mu},\mathbf{B})=\mathcal{K}_{t}(\mathbf{A},\mathbf{B})$ for any scalar $\mu$, Equation˜10 can be used to compute the block-CG iterate for multiple values of $\mu$ using the same $\mathbf{Q}_{t}$ and $\mathbf{T}_{t}$. For more details on block-Lanczos and the connection to block-CG, see e.g., [54, 62, 8, 67].
There are many variants of block-Lanczos, since there are many possible orthogonlization schemes that can be used within block-Lanczos type algorithms [7, 53, 1]. The following provides a high-level overview of the costs associated with block-Lanczos with full-reorthogonalization.
@@
  • $O(dm^{2}t^{2})$ floating point operations (in addition to products with $\mathbf{A}$), and
  • $O(dmt)$ storage (in addition to storage required for $\mathbf{A}$).
@@
Proof 3.8.
Let $n=\dim(\mathcal{K}_{t}(\mathbf{A},\mathbf{B}))$. Since $\mathbf{B}$ has $m$ columns, by the definition Equation˜6 of the block Krylov subspace, $n\leq\min\{d,mt\}$. At each iteration the block-Lanczos algorithm performs one matrix-load ($m$ parallel matrix products) and lower-order arithmetic. The dominant storage cost is storing $\mathbf{Q}_{t}$, which has $n\leq mt$ columns of length $n$. The arithmetic cost is dominated by the $O(dn^{2})=O(dm^{2}t^{2})$ operations required to obtain $\mathbf{Q}_{t}$ using full reorthogonalization. Recall that $\mathbf{T}_{t}$ (and hence $\mathbf{T}_{t}+\mu\mathbf{I}$) is a $n\times n$ matrix with bandwidth $O(m)$. Thus, the linear system $(\mathbf{T}_{t}+\mu\mathbf{I})\mathbf{z}=\mathbf{e}_{1}$ can be solved in $O(nm^{2})=O(m^{3}t)$ time. Subsequently, since $\mathbf{Q}_{t}$ is a $d\times n$ matrix, $\mathbf{Q}_{t}\mathbf{z}$ can be computed in $O(dn)=O(dmt)$ time.
We emphasize that, similar to unpreconditioned CG (see e.g., [39, 28, 5]), after the block-Lanczos algorithm has been run, the cost to compute the block-CG iterates for multiple values of $\mu$ is relatively small. This is in contrast to Nyström PCG, which is is not particularly well suited for solving Equation˜1 for multiple values of $\mu$; while the Nyström approximation can be reused, PCG must be re-run and new products with $\mathbf{A}$ computed.
Notice that the costs (besides matrix-loads) scale with the block-size $m$. We again remind the reader that, while Theorem˜3.4 guarantees augmented block-CG outperforms Nyström preconditioning in terms of matrix-loads, our theory provides no guarantees for other cost matrices like matrix-products or floating point operations. Users should carefully consider the costs and benefits associated with the various algorithms in the context of their own computing environments.
@@
By avoiding orthogonalizing and storing the whole Krylov basis, it is in fact possible to reduce the floating point cost to $O(dmtL+m^{3}tL)$ and the storage to $O(dm+dL)$, where $L$ is the number of values of $\mu$ at which one wishes to evaluate $\mathsf{bcg}_{t}^{(1)}(\mu)$ [54]. However, there are several caveats to such an implementation. First, such methods require knowing the values of $\mu$ at which one wishes to evaluate $\mathsf{bcg}_{t}^{(1)}(\mu)$ ahead of time. Second, the finite-precision behavior of such methods can differ greatly from the exact arithmetic behavior. We explore the impacts of finite precision arithmetic ˜A, where we perform some numerical experiments that indicate that full reorthogonalization may be unnecessary in some situations.
For methods based on the standard Lanczos algorithm, there is quite a bit of theory which guarantees that such methods can still work, even without orthogonalization [55, 56, 20, 32, 52, 8]. Block Krylov methods must contend with additional difficulties such as blocks becoming ill-conditioned or even rank-deficient as the iteration proceeds (which can happen even in exact arithmetic). Unfortunately, much less theory is known about the block-Lanczos algorithm in finite precision arithmetic [75]. Without reorthogonalization, we observe that in many cases the block-CG method fails to converge at all, while in other cases it does converge for some time; see ˜A. Understanding this behavior is well beyond the scope of the current work.